OpenStudy (anonymous):

find the slope of the line that passes through the pair of points (6.5, 10.5) and (–9.5, 7)

OpenStudy (anonymous):

i dont get what im doing wrong i followed the last steps you did

OpenStudy (blurbendy):

(7 - 10.5) = -2.5 -9.5 - 6.5 = 16 -2.5 / 16 = -5 / 32 slope = -5/32 Use the formula (Y2 - Y1) / (X2 - X1) Remember, two negatives make a posi.tive

OpenStudy (anonymous):

my answer is is supposed to be one of the following a. 4.57 c. 40 b. 0.22 d. –1.14

OpenStudy (anonymous):

how do i change that to a decimal

OpenStudy (anonymous):

5/32 ?

OpenStudy (blurbendy):

10.5 - 7 should be 3.5 not 2.5 sorry

OpenStudy (blurbendy):

so it's 3.5 / 16 = ~ 0.22

OpenStudy (anonymous):

thank you for showing me all of your work i am just coming back to school and it has been really hard for me to remember formulas and how do do things, do you recommend and good refresher books?

OpenStudy (anonymous):

awesome ill have to check them out. thank you.

OpenStudy (blurbendy):

also check out www.khanacademy.com

OpenStudy (blurbendy):

go to the math section. soooooo helpful.